OpenStudy (anonymous):

what is the anti derivative of 2x/4

OpenStudy (anonymous):

Anti derivative means Integral...

OpenStudy (anonymous):

\[\int\limits(\frac{2x}{4}).dx = \frac{2}{4} \int\limits (x).dx\]

OpenStudy (anonymous):

And : \[\int\limits x^n dx = \frac{x^{n+1}}{n+1}\]